Q . THE EQUATION ax2 + 2bxy + by2 =0 REPRESENTS A PAIR OF LINES. COMBINED EQUATION OF LINES THAT CAN BE OBTAINED BY REFLECTING THESE LINES ABOUT X-AXIS IS -
  1. bx2 – 2bxy + ay2 = 0
  2. ax2 + 2bxy +by2 = 0
  3. bx2 + 2bxy + ay2 =0
  4. ax2 – 2bxy + by=0

Grade:Select Grade

2 Answers

Himaja
9 years ago
Let (X,Y) be a point in the req locus.
Clearly (X,Y) is the image of the point (X,-Y) wrt y=0.
According to the question, (X,-Y) lies on ax2+2bxy+by2=0
Therefore aX2- 2bXY +bY2= 0
(X,Y) an arbitrary point in the locus satisfyies the above equation the req locus is ax2-2bxy+by2=0
Alternate:
To find equation of a curve obtained by reflecting the given curve about x axis just replace (x,y) with( x.-y) you get the eqn of req curve.
If reflection is taken about y axis then replace (x,y) with (-x,y)
If you like my answer pls approve it.
Bharat Makkar
9 years ago
thanks.......